A sample of oxygen at 44◦C occupies 373 mL.

If this sample later occupies 1102 mL at 60◦C

and 1.8 atm, what was its original pressure?

Answer in units of atm


1
Expert's answer
2022-05-16T10:50:06-0400

T1=44 °C=317 K

V1=373 ml=0.000373 m3

T2=60°C=333 K

V2=1102 ml=0.001102 m3

P2=1.8 atm=182385 Pa

R=8.31 J/(K mole)

P1 -?

"PV=nRT"

"n=\frac{P_2V_2}{RT_2}=\frac{182385x0.001102}{8.31x333}=0.0726 mole"

"P_1=\frac{nRT_1}{V_1}=\frac{0.0726x8.31x317}{0.000373}=512729 Pa"